OpenStudy (anonymous):

60 gal tank of water is leaking, with 57 gal left after 10 min, 54 gal left after 20 min. Find the slope of the line/ write the linear equation for # of gal left in tank after x amount of min.

hero (hero):

Hint: \[(x_1, y_1) = (10, 57)\]\[(x_2,y_2) = (20,54)\] Find the slope m, using the slope formula Then use the point slope formula: \[y - y_1 = m(x - x_1)\] Afterwards, place the equation of the line in the form y = mx + b

OpenStudy (anonymous):

slope is -.3?

hero (hero):

How did you get that?

OpenStudy (anonymous):

54-57/20-10

hero (hero):

I see. Okay. People normally write the slope in the form of a fraction. Good job so far.

OpenStudy (anonymous):

ok, so Y=-3/10 (10) +b?

OpenStudy (anonymous):

can you write out the equation? I am getting confused..

hero (hero):

You have to input the y value as well along with the x value if you want to find b

OpenStudy (anonymous):

y-57=-3/10(x-10)?

hero (hero):

You could do it that way as well. I'm used to using the point slope formula so I would do it that way, however, doing it your way may be easier. If you begin with y = mx+b and you have found the slope, then you could input the point and the slope into that formula and then solve for b: We know that the slope is m = -.3 and one of the points is (x,y) = (10,57). So to find the b value, y = mx + b becomes: 57 = .3(10) + b Can you finish up here by solving for b?

hero (hero):

No matter which method you use, you should end up with the same answer in the form y = mx + b. Using your method, once you find b, you would rewrite the slope-intercept formula, but include the values for m and b only.

OpenStudy (anonymous):

I think I may have forgotten a rule here. Why does -.3 become .3?